OpenStudy (anonymous):

what is the sum of 3 1/3 and 4 2/7?

OpenStudy (anonymous):

1+8/7 15/7

OpenStudy (anonymous):

fine?

OpenStudy (anonymous):

thats not in the answer choice.

OpenStudy (anonymous):

but is 3(1/3) and 4(2/7) ok?

OpenStudy (anonymous):

rewrite it as: 10/3 + 30/7 common denominator is 21 so you mutiply 10/3 x 7/7 and 30/7 x 3/3 you get 70/21 + 90/21 = 160/21

OpenStudy (anonymous):

A.7 3/21 B.7 3/10 C.7 3/7 D.7 13/21 And yes im sorry i wasnt meaning to rush you I thought u was saying the answer

OpenStudy (ragingsquirrel):

D is the answer.

OpenStudy (anonymous):

3 1/3 and 4 2/7 = 3(1/3)+4(2/7) = 3(7/21)+4(6/21)

OpenStudy (anonymous):

fine?